OpenStudy (anonymous):

Solve for x, given the equation Square root of x minus 5 + 7 = 11.

OpenStudy (anonymous):

\[\sqrt{x}-5+7 = 11 \] or \[\sqrt{x-5}+7 = 11?\]

OpenStudy (anonymous):

or is it all under sqaure root?

OpenStudy (anonymous):

if someone could show mw thw general process or pattern of how to do this equations and others similar that would be helpful.

OpenStudy (anonymous):

\[\sqrt{x}-5+7=11\]

OpenStudy (anonymous):

Add 5 and subtract 7 to both sides \[\sqrt{x} = 9\]

OpenStudy (anonymous):

81 = x (9)^2 = 81

OpenStudy (anonymous):

x-5 is square rooted

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sqrt{x-5}+7 = 11\]

OpenStudy (anonymous):

i got x=11.. :(

OpenStudy (anonymous):

subtract 7 from both sides sqrt(x-5) = 4 sqaure both sides x - 5 = 16 +5 +5 x = 21

OpenStudy (anonymous):

how do i know if its extreneous or not ?

OpenStudy (anonymous):

wow im really not getting this.

OpenStudy (anonymous):

oh wait i just squared both sides. ok

OpenStudy (anonymous):

The value under the sqaure root is not negative.

OpenStudy (anonymous):

what did you do after you got x=21?

OpenStudy (whpalmer4):

\[\sqrt{x-5}+7=11\]Move the 7 to the right by subtracting it from both sides: \[\sqrt{x-5}=4\]Square both sides \[(x-5) = \pm 16\]Solve for both values of \(x\) \[x-5=+16\]\[x-5=-16\]

OpenStudy (whpalmer4):

Now, here's the crucial part: whenever you solve by squaring a square root like this, you MUST check to see if all of your solutions actually work in the original equation! Many times, one of them will not, and such solutions are called extraneous solutions.

OpenStudy (anonymous):

ok but now im at square root thingy over 16 then +7=11. what to do now?

OpenStudy (whpalmer4):

Let's test \(x = 21\) \[\sqrt{21-5} + 7 = 11\]\[\sqrt{16}+7 = 11\]\[4+7=11\checkmark\]That one is good. However, we also got \[x-5=-16\]\[x = -16+5 = -11\] \[\sqrt{-11-5}+7 = 11\]\[\sqrt{-16} + 7 = 11\]\[4i + 7 = 11\]That's not true, so that is an extraneous solution.

OpenStudy (whpalmer4):

\[\sqrt{16} = \pm4\]

OpenStudy (anonymous):

oh i see

OpenStudy (whpalmer4):

because 4*4 = 16 and (-4)*(-4) = 16

OpenStudy (anonymous):

ok,so how do i get 16 out of square root?

OpenStudy (whpalmer4):

If you didn't know this, you could factor 16: 16 = 2*2*2*2 so \[\sqrt{16} = \sqrt{2*2*2*2} = \sqrt{(2*2)*(2*2)} = 2*2 = 4\]because \[\sqrt{a*a} = a\]so long as \(a\ge0\)

OpenStudy (anonymous):

hm ok. i see that.but is 4 going to then be divided into 16?

OpenStudy (anonymous):

like where do i go from sqaure root over 16=4? i know 4 goes into it but

OpenStudy (whpalmer4):

there's a convention that the square root sign means the positive or principal square root, so even though \((-4)*(-4) = 16\), \(\sqrt{16} = 4\). When we square both sides, we "forget" whether that was the positive or negative value, so we have to stick in the \(\pm\). \[\sqrt{16} = 4\] I'm not sure what you're asking...

OpenStudy (anonymous):

im asking, im not sure

OpenStudy (anonymous):

so i guess that is the answer

OpenStudy (whpalmer4):

what is the answer?

OpenStudy (anonymous):

what do i do if there is a number i front of a square root?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

i really appreciate this help

OpenStudy (whpalmer4):

\[i = \sqrt{-1}\]so \[i^2 = -1\] Yes, \(x=21\) is the answer to this problem. We can test it: \[\sqrt{x-5}+7=11\]\[\sqrt{21-5}+7=11\]\[\sqrt{16}+7=11\]\[4+7=11\]\[11=11\]

OpenStudy (whpalmer4):

Don't be too concerned if the \(i\) stuff confuses you for a while — it took mathematicians several centuries to figure it out the first time :-)

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

well im really thanking them now.. geeze

OpenStudy (whpalmer4):

in fact, just negative numbers were regarded as highly suspicious for many years!

OpenStudy (anonymous):

well what do i do if i have an equation similar to this one but there is a nnumber on the outside of the square root

OpenStudy (anonymous):

lol

OpenStudy (whpalmer4):

Not a problem, let's do one! \[2\sqrt{x-5} + 2 = 10\]

OpenStudy (anonymous):

oh goodie

OpenStudy (whpalmer4):

Start the same way, shovel all the loose numbers over to the other side: \[2\sqrt{x-5}=8\]Now we have a choice, we can divide through by 2, giving us something like we had before, or we can just keep it there. Let's keep it, and square both sides: \[(2\sqrt{x-5})^2 = 64\]

OpenStudy (anonymous):

hmm ok

OpenStudy (anonymous):

im follpowiing

OpenStudy (whpalmer4):

we can expand that to \[2*2*(\sqrt{x-5})^2 =\pm 64\]Divide both sides by 4\[(\sqrt{x-5})^2 = \pm16\] squaring the square root just gives us the stuff under the square root sign (but don't forget about the need for the \(\pm\) and the extraneous solution check)

OpenStudy (whpalmer4):

so that gives us\[x-5 =\pm16\] and \[x = 21, x = -11\]just like the previous problem :-) Still, we need to check for extraneous solutions: \[2\sqrt{21-5}+2 = 10\]\[2\sqrt{16}+2=10\]\[2*4+2=10\checkmark\] \[x=21\] is a good solution \[2\sqrt{-11-5}+2 = 10\]\[2\sqrt{-16}+2=10\]\[2*4i+2\ne10\]And just like before \(x=-11\) is an extraneous solution.

OpenStudy (whpalmer4):

Make sense?

OpenStudy (anonymous):

ok owd you get x-5-=16

OpenStudy (anonymous):

:o

OpenStudy (whpalmer4):

\[(\sqrt{x-5})^2 = \pm16\] If we square the square root of something, we just get the something: \[(\sqrt{a})^2 = \sqrt{a}\sqrt{a} = \sqrt{a*a} = a\]

OpenStudy (anonymous):

hmm